A bar AB is hinged to the foundation at A and supported by strut CD is subjected to a horizontal 50 kN load at B,as shown in fig Determine the nature and magnitude of the forces in strut and also the reaction at A

1
Expert's answer
2022-01-26T04:36:02-0500


SOKN sino Solo =/0=33.71 tos (p= 09 1,83 1. 2.2.0 10.00 c 11.00 4.22m F 2- 12.00 D 183 m 1.93m 2.2 LUNCH t and = 1.22 0-ton, → SOKN (SOKN cos FAxial force in the stunt about Ai 150 KN sind) 4y = F 44 F = 100kN sind F= 200kN (1.22" > F= 110.9 KN And. AB= (3-66) + (2.44)² = 4.4m Force along 50 KN coso = N N = SOKN (18) A/N = 41.6 kN N horizontal EN CASO = 41.6 KN * 11,83 = 34,6 KN IN x = 34.6 KN Nsind = 41.6 kN x = 23,07 kn Ny = 23,07 KN 02.00 the rod, 03.00 04.00 - 05.00 Nvertical Eve A 2.2